OpenStudy (anonymous):

Simplify the following expression completely. (18)/(6 x -30)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[\frac{18}{6x-30}\]?

OpenStudy (jhannybean):

yeah.

OpenStudy (anonymous):

you can factor out a \(6\) from the denominator, then cancel

OpenStudy (jhannybean):

factor out a 6 from the denominator

OpenStudy (anonymous):

wait... factor out 6 .. how so?

OpenStudy (jhannybean):

6 x what = 30?

OpenStudy (anonymous):

5

OpenStudy (jhannybean):

what = 30/6 = ?

OpenStudy (jhannybean):

Ok good. so when you factor out a 6, you will have 6(x - 5)

OpenStudy (jhannybean):

Do you see how that works?

OpenStudy (anonymous):

so my answer would be 18/ 6(x-5)

OpenStudy (jhannybean):

What is 18/6?

OpenStudy (anonymous):

it will be after you cancel the common factor of 6 top and bottom

OpenStudy (jhannybean):

6 x what = 18?

OpenStudy (anonymous):

3

OpenStudy (jhannybean):

So you will have 3/(x-5)

OpenStudy (anonymous):

okay thanks

OpenStudy (jhannybean):

No problem :)